# Ultraluminous X-ray sources in Globular Clusters

### v1

@@ 3.4 ULX properties
The diversity in donor types across ULX groups highlights the various evolutionary pathways that can lead to ULX formation. MS, HG, and RG donors are all well-represented, suggesting that ULXs can form and persist across a wide range of stellar evolutionary stages
@@
666 Table presents the fractional distribution of ULXs across different groups for various simulations. For group definitions see Table 3
Table 4 shows a modest anticorrelation between the relative size of the $G_{\textrm{IMBH}}$ and BH ULXs (Kendall [^7] Kendall’s $\tau$ is preferable to Pearson’s $r$ when working with small samples and non-normal distributions, while still providing a similarly interpretable measure of association between $-1$ and $+1$.’s $\tau=-0.49,-0.38,-0.37$, for $G_{\textrm{BH,MS}}$, $G_{\textrm{BH,HG}}$, and $G_{\textrm{BH,CHeB}}$, respectively). The presence of IMBH results in a continuous formation of IMBH powered ULX though the cluster history which, being an independent sourse to other BH ULX groups present mostly in the early cluster evolution phases, lowers their fractional imput to the ULX population. IMBH quickly removes BHs and possibly NSs from the system (e.g. Hong et al. 2020).
3.4.1 IMBH ULXs
In this study, we define intermediate-mass black holes (IMBHs) as those with masses exceeding $500\mathrel{\mathrm{M}_{\odot}}$. IMBH ULXs were observed in 6 simulations, all featuring multiple stellar populations (npop2) and non-tidally filling initial conditions (nTF). When present, IMBHs contribute significantly to the ULX population, accounting for 16-44% of ULXs over a Hubble time (Table 4).
IMBHs form rapidly, within several Myr after the zero-age main sequence (ZAMS), classifying them as ”fast IMBHs” (Giersz et al. 2015). In simulations with delayed second populations, IMBH formation occurs 100+ Myr after the initial population. Notably, IMBH ULXs form exclusively in non-tidally filling (nTF) simulations, strenghtening the preference for denser stellar environments.
Table 5: IMBH ULX properties
label $\mathrm{t}_{\mathrm{phys,min}}$ $\mathrm{t}_{\mathrm{phys,max}}$ duty-cycle [%] companions
npop2-cpop05-fb10-nTF $7.95$ $13,306.24$ $0.06$ MS: 22, HeWD: 4, COWD: 3, TPAGB: 2, CHeB: 2
npop2-cpop05-fb10-nTF-NF $100.63$ $14,925.72$ $0.02$ MS: 41, COWD: 5, RG: 3, HeWD: 2, CHeB: 1,
EAGB: 1, TPAGB: 1, ONeWD: 1
npop2-cpop05-fb95-nTF $11.49$ $13,791.13$ $0.87$ MS: 29, COWD: 13, HG: 1, HeWD: 1, RG: 1
npop2-cpop05-fb95-nTF-NF $104.89$ $14,751.19$ $0.22$ MS: 89, COWD: 12, TPAGB: 2, HeWD: 2, CHeB: 2,
HeGB: 1, ONeWD: 1, HG: 1
npop2-cpop2-fb10-nTF $9.19$ $14,171.06$ $0.14$ MS: 32, COWD: 5, HeWD: 2, RG: 1, TPAGB: 1,
ONeWD: 1
npop2-cpop2-fb95-nTF $8.99$ $14,546.35$ $0.43$ MS: 43, COWD: 11, HG: 3, HeWD: 3, TPAGB: 2,
RG: 2
888 Table presents the properties of $6$ IMBH which appeared in our simulations.
[…]
IMBH ULXs formed exclusively in simulations with two stellar populations and non-tidally filling initial conditions. The centrally concentrated second population facilitates stellar interactions and mergers, while nTF clusters retain more stars, particularly black holes, in central regions, extending the IMBH formation window.
3.5 ULX progenitors
Table 6: Initial Properties of ULX Progenitors by Group
999 Values are provided as median with 95% confidence intervals or averages. Groups are defined in Table 3 and accompaning text. * denotes component (A - accretor, D - donor). $\mathrm{M}_{\mathrm{ZAMS}}$: ZAMS mass ($\mathrel{\mathrm{M}_{\odot}}$); esc: fraction of ULXs formed from escapers; bin: binary fraction among progenitors; mode: dominant formation mode (see Table 2); $\mathrm{f}_{\mathrm{prim}}$: fraction of progenitors formed from primary stars; $\mathrm{f}_{\mathrm{pris}}$: fraction of pristine ULXs formed; $\mathrm{M}_{\mathrm{comp}}$: companion ZAMS mass ($\mathrel{\mathrm{M}_{\odot}}$); $a$: ZAMS separation ($\mathrel{\mathrm{R}_{\odot}}$); $e$: ZAMS eccentricity. Mode, $\mathrm{f}_{\mathrm{prim}}$, $\mathrm{f}_{\mathrm{pris}}$, $\mathrm{M}_{\mathrm{comp}}$, $a$, and $e$ values are calculated only for binary progenitors.
3.6 ULX descendants
Table 7: ULX descendants
D E X M O
npop1-fb10-TF $67$ $14$
npop1-fb10-nTF $4$ $22$ $46$ $25$ $4$
npop1-fb95-TF $222$ $36$ $1$
npop1-fb95-nTF $18$ $69$ $152$ $70$ $19$
npop2-cpop05-fb10-TF $2$ $35$ $11$ $13$
npop2-cpop05-fb10-TF-NF $2$ $25$ $21$ $14$ $2$
npop2-cpop05-fb10-nTF $18$ $20$ $10$ $42$ $1$
npop2-cpop05-fb10-nTF-NF $23$ $12$ $10$ $65$ $1$
npop2-cpop05-fb95-TF $4$ $117$ $30$ $37$ $3$
npop2-cpop05-fb95-TF-NF $13$ $90$ $49$ $37$ $9$
npop2-cpop05-fb95-nTF $55$ $72$ $38$ $81$ $3$
npop2-cpop05-fb95-nTF-NF $66$ $54$ $41$ $136$ $8$
npop2-cpop2-fb10-TF $37$ $2$ $12$ $3$
npop2-cpop2-fb10-TF-NF $3$ $24$ $15$ $14$ $1$
npop2-cpop2-fb10-nTF $17$ $14$ $16$ $55$ $2$
npop2-cpop2-fb10-nTF-NF $9$ $27$ $60$ $67$ $12$
npop2-cpop2-fb95-TF $2$ $115$ $17$ $34$ $9$
npop2-cpop2-fb95-TF-NF $4$ $88$ $42$ $36$ $13$
npop2-cpop2-fb95-nTF $73$ $41$ $105$ $66$
npop2-cpop2-fb95-nTF-NF $17$ $60$ $124$ $130$ $45$
The fate of ULX systems after their active phase provides valuable insights into their evolutionary pathways and potential contributions to other astrophysical phenomena. Table 7 summarizes the various outcomes for ULX systems across different simulation configurations.
ULX systems can undergo several fates: disruption, escape from the cluster, exchange interactions, mergers, or other outcomes. The distribution of these fates appears largely stochastic across different simulation parameters. Systems categorized as ”other” typically experience no significant events for the remainder of their evolution and are predominantly NS/BH + WD or DCOs at the end of simulation.
Merger events typically occur $0.79^{+0.74}_{-0.53}\mathrel{\mathrm{Myr}}$ after the ULX phase. Double compact objects (DCOs) form in $\sim 63\%$ of cases, with $\sim 34\%$ of them being ejected from the cluster before merger or the end of simulation. These DCOs that remain in the cluster may contribute to the population of gravitational wave (GW) sources with merger rate of $\sim 11\%$.
3.7 ULXs among escapers
Table 8: Properties of Escaper and In-cluster ULXs
escaper in-cluster
NS BH NS BH
Typ. companion HeMS (70%) CHeB (80%) HG (74%) HG (55%)
Escaper ULXs exhibit several distinct characteristics compared to their in-cluster counterparts. They tend to have lower-mass accretors and donors, with median masses of $13.00$ and $6.05\mathrel{\mathrm{M}_{\odot}}$, respectively, compared to $22.24$ and $39.34\mathrel{\mathrm{M}_{\odot}}$ for in-cluster ULXs. Escaper ULXs also have tighter orbits (median $57.42$ vs. $119.93\mathrel{\mathrm{R}_{\odot}}$) and a significantly higher fraction of neutron star accretors ($40\%$ vs. $4\%$). The median maximum X-ray luminosity of escaper ULXs ($\log_{10}\mathrm{L}_{\mathrm{X,max}}=39.89\mathrel{\mathrm{erg}/\mathrm{s}}$) is slightly lower than that of in-cluster ULXs ($\log_{10}\mathrm{L}_{\mathrm{X,max}}=40.26\mathrel{\mathrm{erg}/\mathrm{s}}$).
The ejection processes preferentially remove less massive systems, explaining the lower masses observed in escaper ULXs. The higher fraction of neutron star accretors among escapers could be due to the higher natal kicks and retention of more massive black holes within the cluster potential.
These findings have important implications for understanding the origin and evolution of field ULXs. A significant fraction of observed field ULXs may have originated in globular clusters, which could explain some of the observed diversity in ULX populations. Furthermore, the differences between escaper and in-cluster ULXs provide valuable insights into how environment and dynamical history can influence the properties of these extreme systems.
4 Discussion
A key limitation in the broader study of ULXs in GCs is the paucity of confirmed observations within the Milky Way and its immediate vicinity. The substantial distances involved often complicate the detection and characterization of low-luminosity donor stars, which are believed to comprise a significant fraction of ULX donors (e.g. Wiktorowicz et al. 2017).
4.1 Comparison with field populations
Our escaper NS ULXs have a median maximum X-ray luminosity of $\log_{10}\mathrm{L}_{\mathrm{X,max}}=39.22\mathrel{\mathrm{erg}/\mathrm{s}}$, with a few BH ULXs exceeding $10^{40}\mathrel{\mathrm{erg}/\mathrm{s}}$. This stays in contrast to the field populations, where a significant fraction of more luminous sources is also expected (e.g., Wiktorowicz et al. 2015). In the case of globular clusters, we find that such extreme ULXs are more likely to be retained within the cluster.
[…]
4.2 Counterparts
The identification and characterization of ULX counterparts is crucial for understanding the nature of these systems, yet it remains a significant observational challenge (Heida et al. 2016; Wiktorowicz et al. 2021). While many ULXs are found in star-forming regions, their faintness in the optical and near-infrared (NIR), combined with the crowded environments, makes it difficult to confirm donor stars (López et al. 2020; Allak et al. 2022; Heida et al. 2019). The observed optical emission can be a combination of light from the accretion disc and/or the donor star (Heida et al. 2016). Accurate astrometry is essential for identifying counterparts, particularly for distinguishing between potential candidates within the error radii of ULX positions (Allak et al. 2022).
Red supergiants (RSGs) have also been identified as potential donors in ULXs (Heida et al. 2016, 2019; López et al. 2020). However, several sources initially classified as RSGs have been found to be nebulae, making the classification challenging (López et al. 2020). Actually, the fraction of ULXs with RSG donors is potentially higher than predicted by binary evolution models (López et al. 2020).
It’s important to note that observational biases exist that favour the detection of more luminous and massive companions, such as RSGs (e.g. Heida et al. 2019). Additionally, systems may be old, which makes it less likely for the donor to be a supergiant (Wiktorowicz et al. 2017). Therefore, ULX counterparts remain ’candidates’ and need more thorough survey searches (Wiktorowicz et al. 2021).
4.3 Expectations for Antennae galaxy
The correlation between X-ray source positions and stellar clusters has been previously noted in starburst galaxies (e.g. Kaaret et al. 2004). The study of the Antennae galaxies by Poutanen et al. (2013) reveals a highly significant association between bright X-ray sources and stellar clusters, although most of these X-ray sources are located outside of the clusters.
This finding is consistent with the idea that many ULXs are massive X-ray binaries that have been ejected from their birth clusters, rather than being intermediate mass black holes (IMBHs). Both ULXs and XRBs show a strong association with young stellar clusters, indicating a common origin.
The displacements of ULXs from centres of their neighbouring star clusters are statistically significant and range up to 300 parsecs. This is likely due to the ejection of massive binaries from the clusters.
The ejection mechanism suggested by Poutanen et al. (2013), supernova kicks and few-body encounters are consistent with the results of our work, where ULX among escapers have much higher NS accretor fraction and much smaller separations.
Poutanen et al. (2013) provides evidences that dynamical ejection can be a major formation mechanism for these systems in the field.
4.4 Future prospects
Future studies of ULXs in globular clusters should include more simulations to better estimate statistical effects, errors, and the significance of conclusions. A more consistent survey of the parameter space is needed, as well as the incorporation of evolutionary parameters (both stellar and binary evolution) beyond dynamical ones.
In the future work we plan to include the effect of beaming in ULXs (e.g. Lasota & King 2023). Beaming increases the apparent luminosity of ULXs, but decreases the probability of observation due to possible misalignment (Wiktorowicz et al. 2019; Khan et al. 2022). Population synthesis simulations, Wiktorowicz et al. (e.g. 2019) indicate that the majority of neutron star (NS) ULXs are beamed. The beaming factor is dependent on the mass transfer rate, with higher mass transfer rates resulting in stronger beaming (King 2009).
Additionally, the importance of wind Roche-lobe overflow (WRLOF; see e.g., Wiktorowicz et al. 2021; Zuo et al. 2021, and references therein) for NS ULXs with (super)giant donors should be explored. WRLOF can lead to stable mass transfer even with large mass ratios, and can boost mass transfer rates to reach ULX luminosity levels.
5 Conclusions
The apparent absence of ULXs in Milky Way GCs aligns more closely with our models of initially tidally filling clusters rather than non-tidally filling ones.
Furthermore, our results suggest that field populations may be significantly ”polluted” by ULXs ejected from GCs (escapers), which could contribute to the relative underrepresentation of ULXs in GCs compared to field environments. Escapers can have properties very similar to the field populations or form dinsinct configurations.
Acknowledgments
GW, MG, AH, LH were supported by the Polish National Science Center (NCN) through the grant 2021/41/B/ST9/01191 AA acknowledges support for this paper from project No. 2021/43/P/ST9/03167 co-funded by the Polish National Science Center (NCN) and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339. For the purpose of Open Access, the authors have applied for a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
Data Availability
Data available on request.

### v3

@@ 3.4 ULX properties
The diversity in donor types across ULX groups highlights the various evolutionary pathways that can lead to ULX formation. Donors of types MS, HG, and RG are all well-represented, suggesting that ULXs can form and persist across a wide range of stellar evolutionary stages.
Table 8 shows a modest anticorrelation between the relative size of the $G_{\textrm{IMBH}}$ and BH ULXs (Kendall [^8] Kendall’s $\tau$ is preferable to Pearson’s $r$ when working with small samples and non-normal distributions, while still providing a similarly interpretable measure of association between $-1$ and $+1$.’s $\tau=-0.49,-0.38,-0.37$, for $G_{\textrm{BH,MS}}$, $G_{\textrm{BH,HG}}$, and $G_{\textrm{BH,CHeB}}$, respectively). The presence of IMBH results in a continuous formation of IMBH powered ULX though the cluster history which, being an independent source to other BH ULX groups present mostly in the early cluster evolution phases, lowers their fractional input to the ULX population. IMBH quickly removes BHs and possibly NSs from the system (e.g., Hong et al., 2020).
3.5 IMBH ULXs
In this study, we define IMBHs as those with masses exceeding $500\mathrel{\mathrm{M}_{\odot}}$. IMBH ULXs were observed in 6 simulations, all featuring multiple stellar populations (npop2) and non-tidally filling initial conditions (nTF). When present, IMBHs contribute significantly to the ULX population, accounting for 16-44% of ULXs over a Hubble time (Table 8).
IMBHs form rapidly, within several Myr after the ZAMS, classifying them as ”fast IMBHs” (Giersz et al., 2015). In simulations with delayed second populations, IMBH formation occurs 100+ Myr after the initial population. Notably, IMBH ULXs form exclusively in non-tidally filling (nTF) simulations, strenghtening the preference for denser stellar environments.
Table 3: IMBH ULX properties
label $\mathrm{t}_{\mathrm{phys,min}}$ $\mathrm{t}_{\mathrm{phys,max}}$ $\text{M}_{\mathrm{min}}$ $\text{M}_{\mathrm{max}}$ duty-cycle [%] companions
npop2-cpop05-fb10-nTF $0.008$ $13.3$ $3.8$ $13.0$ $0.06$ MS: 22, HeWD: 4, COWD: 3,
TPAGB: 2, CHeB: 2
npop2-cpop05-fb10-nTF-NF $0.1$ $14.9$ $0.5$ $13.5$ $0.02$ MS: 41, COWD: 5, RG: 3,
HeWD: 2, CHeB: 1, EAGB: 1,
TPAGB: 1, ONeWD: 1
npop2-cpop05-fb95-nTF $0.01$ $13.8$ $5.0$ $13.5$ $0.87$ MS: 29, COWD: 13, HG: 1,
HeWD: 1, RG: 1
npop2-cpop05-fb95-nTF-NF $0.1$ $14.8$ $0.6$ $12.9$ $0.22$ MS: 89, COWD: 12, TPAGB: 2,
HeWD: 2, CHeB: 2, HeGB: 1,
ONeWD: 1, HG: 1
npop2-cpop2-fb10-nTF $0.009$ $14.2$ $0.6$ $4.5$ $0.14$ MS: 32, COWD: 5, HeWD: 2,
RG: 1, TPAGB: 1, ONeWD: 1
npop2-cpop2-fb95-nTF $0.009$ $14.5$ $2.1$ $6.8$ $0.43$ MS: 43, COWD: 11, HG: 3,
HeWD: 3, TPAGB: 2, RG: 2
999 Properties of $6$ intermediate-mass black holes (IMBHs) ULXs observed in our simulations. Each IMBH forms multiple, short-lived ULX phases with different companions. Column definitions: label - simulation model name; $\mathrm{t}_{\mathrm{phys,min}}$ minimum ULX phase onset age ($\mathrel{\mathrm{Gyr}}$); $\mathrm{t}_{\mathrm{phys,max}}$ maximum age at which ULX phases occur ($\mathrel{\mathrm{Gyr}}$); $\text{M}_{\mathrm{min}}$ minimum IMBH mass during ULX phases ($1000\;\times\mathrel{\mathrm{M}_{\odot}}$); $\text{M}_{\mathrm{max}}$ maximum IMBH mass during ULX phases ($1000\;\times\mathrel{\mathrm{M}_{\odot}}$); duty-cycle - percentage of total IMBH lifetime spent in ULX phases; companions - types and numbers of donor stars (MS - main sequence, HG - Hertzsprung gap, RG - red giant, CHeB - core helium burning, EAGB - early asymptotic giant branch, TPAGB - thermally pulsating asymptotic giant branch, HeGB - helium giant branch, HeWD - helium white dwarf, COWD - carbon-oxygen white dwarf, ONeWD - oxygen-neon white dwarf).
[…]
IMBH ULXs formed exclusively in simulations with two stellar populations and non-tidally filling initial conditions. The centrally concentrated second population facilitates stellar interactions and mergers, while nTF clusters retain more stars, particularly BHs, in central regions, extending the IMBH formation window.
IMBH ULXs typically exist throughout the entire GC life, with first ULX phases appearing as early as $\sim 8\mathrel{\mathrm{Myr}}$ in some models and continuing until the end of our simulations at $\sim 14\mathrel{\mathrm{Gyr}}$, indicating that IMBH ULXs can form very late in GC evolution.
3.6 ULX progenitors
Table 4: Initial properties of ULX progenitors by group
101010 Initial properties of ULX progenitors, including median values with $95\%$ confidence intervals or averages. Groups are defined in Table 2 and accompaning text. * denotes component (A - accretor, D - donor). $\mathrm{M}_{\mathrm{ZAMS}}$: ZAMS mass ($\mathrel{\mathrm{M}_{\odot}}$); esc: fraction of ULXs formed from escapers; bin: binary fraction among progenitors; mode: dominant formation mode (see Table 1); $\mathrm{f}_{\mathrm{prim}}$: fraction of progenitors formed from primary stars; $\mathrm{f}_{\mathrm{pris}}$: fraction of pristine ULXs formed; $\mathrm{M}_{\mathrm{comp}}$: companion ZAMS mass ($\mathrel{\mathrm{M}_{\odot}}$); $a$: ZAMS separation ($\mathrel{\mathrm{R}_{\odot}}$); $e$: ZAMS eccentricity. Mode, $\mathrm{f}_{\mathrm{prim}}$, $\mathrm{f}_{\mathrm{pris}}$, $\mathrm{M}_{\mathrm{comp}}$, $a$, and $e$ values are calculated only for binary progenitors.
3.7 ULX descendants
Table 5: ULX descendants
D E X M O
npop1-fb10-TF $67$ $14$
npop1-fb10-nTF $4$ $22$ $46$ $25$ $4$
npop1-fb95-TF $222$ $36$ $1$
npop1-fb95-nTF $18$ $69$ $152$ $70$ $19$
npop2-cpop05-fb10-TF $2$ $35$ $11$ $13$
npop2-cpop05-fb10-TF-NF $2$ $25$ $21$ $14$ $2$
npop2-cpop05-fb10-nTF $18$ $20$ $10$ $42$ $1$
npop2-cpop05-fb10-nTF-NF $23$ $12$ $10$ $65$ $1$
npop2-cpop05-fb95-TF $4$ $117$ $30$ $37$ $3$
npop2-cpop05-fb95-TF-NF $13$ $90$ $49$ $37$ $9$
npop2-cpop05-fb95-nTF $55$ $72$ $38$ $81$ $3$
npop2-cpop05-fb95-nTF-NF $66$ $54$ $41$ $136$ $8$
npop2-cpop2-fb10-TF $37$ $2$ $12$ $3$
npop2-cpop2-fb10-TF-NF $3$ $24$ $15$ $14$ $1$
npop2-cpop2-fb10-nTF $17$ $14$ $16$ $55$ $2$
npop2-cpop2-fb10-nTF-NF $9$ $27$ $60$ $67$ $12$
npop2-cpop2-fb95-TF $2$ $115$ $17$ $34$ $9$
npop2-cpop2-fb95-TF-NF $4$ $88$ $42$ $36$ $13$
npop2-cpop2-fb95-nTF $73$ $41$ $105$ $66$
npop2-cpop2-fb95-nTF-NF $17$ $60$ $124$ $130$ $45$
Approximate Average $19$ $62$ $40$ $51$ $11$
The fate of ULX systems after their active phase provides valuable insights into their evolutionary pathways and potential contributions to other astrophysical phenomena.
Table 5 summarizes the various outcomes for ULX systems across different simulation configurations. The table shows, that ULX systems can undergo several fates: disruption, escape from the cluster, exchange interactions, mergers, or other outcomes. The distribution of these fates appears largely stochastic across different simulation parameters. Systems categorized as ”other” typically experience no significant events for the remainder of their evolution and are predominantly NS/BH + WD or double compact objects at the end of simulation.
Merger events typically occur $0.79^{+0.74}_{-0.53}\mathrel{\mathrm{Myr}}$ after the ULX phase. Double compact objects form in $\sim 63\%$ of cases, with $\sim 34\%$ of them being ejected from the cluster before merger or the end of simulation. These double compact objects that remain in the cluster may contribute to the population of gravitational wave sources with merger rate of $\sim 11\%$.
3.8 ULXs among escapers
Table 6: Properties of escaper and in-cluster ULXs
escaper in-cluster
NS BH NS BH
Typ. companion HeMS (70%) CHeB (80%) HG (74%) HG (55%)
Escaper ULXs exhibit several distinct characteristics compared to their in-cluster counterparts. They tend to have lower-mass accretors and donors, with median masses of $13.00$ and $6.05\mathrel{\mathrm{M}_{\odot}}$, respectively, compared to $22.24$ and $39.34\mathrel{\mathrm{M}_{\odot}}$ for in-cluster ULXs. Escaper ULXs also have a significantly higher fraction of NS accretors ($40\%$ vs. $4\%$) and tighter orbits (median $57.42$ vs. $119.93\mathrel{\mathrm{R}_{\odot}}$). This orbital difference largely stems from dynamical hardening processes that occur prior to ejection, whereby close encounters cause binaries to become more tightly bound (Heggie, 1975). Additionally, the ejection mechanisms preferentially remove harder binaries with sufficient kinetic energy to escape the cluster potential. The median maximum X-ray luminosity of escaper ULXs ($\log_{10}\mathrm{L}_{\mathrm{X,max}}=39.89\mathrel{\mathrm{erg}/\mathrm{s}}$) is slightly lower than that of in-cluster ULXs ($\log_{10}\mathrm{L}_{\mathrm{X,max}}=40.26\mathrel{\mathrm{erg}/\mathrm{s}}$).
The ejection processes preferentially remove less massive systems, explaining the lower masses observed in escaper ULXs. This mass-dependent retention occurs because more massive systems require stronger dynamical encounters to achieve escape velocity from the cluster potential. The higher fraction of NS accretors among escapers could be due to the higher natal kicks and retention of more massive BHs within the cluster potential.
These findings have important implications for understanding the origin and evolution of field ULXs. A significant fraction of observed field ULXs may have originated in GCs, which could explain some of the observed diversity in ULX populations. While our current analysis captures the properties of escaper ULXs at the moment of escape and during the ULX phase, tracking their complete galactic trajectories would require additional post-processing to integrate their orbits through the galactic potential (e.g., Cabrera & Rodriguez, 2023). Furthermore, the differences between escaper and in-cluster ULXs provide valuable insights into how environment and dynamical history can influence the properties of these extreme systems.
4 Discussion
A key limitation in the broader study of ULXs in GCs is the dearth of confirmed observations within the Milky Way and its immediate vicinity. The brightest Milky Way X-ray binaries fall well below the ULX luminosity threshold (Fortin et al., 2023, 2024). The substantial distances involved often complicate the detection and characterization of low-luminosity donor stars, which are believed to comprise a significant fraction of ULX donors (e.g., Wiktorowicz et al., 2017).
4.1 Comparison with field populations
Our escaper NS ULXs have a median maximum X-ray luminosity of $\log_{10}\mathrm{L}_{\mathrm{X,max}}=39.22\mathrel{\mathrm{erg}/\mathrm{s}}$, with a few BH ULXs exceeding $10^{40}\mathrel{\mathrm{erg}/\mathrm{s}}$. This stays in contrast to the field populations, where a significant fraction of more luminous sources is also expected (e.g., Wiktorowicz et al., 2015). In the case of GCs, we find that such extreme ULXs are more likely to be retained within the cluster.
[…]
4.2 Counterparts
Identifying ULX counterparts is crucial yet challenging (Heida et al., 2016; Wiktorowicz et al., 2021). While field ULXs face issues of faintness and crowded environments (López et al., 2020; Allak et al., 2022; Heida et al., 2019), extragalactic GC ULXs present unique challenges due to their unresolved stellar populations. However, GCs’ well-understood stellar demographics suggest these ULXs are likely LMXBs. Rare cases allow indirect donor identification, such as oxygen emission indicating a WD donor in a NGC 4472 GC ULX (Steele et al., 2014). Accurate astrometry remains essential for all ULX studies (Allak et al., 2022), and most counterparts, in both field and GCs, remain tentative due to observational biases (Heida et al., 2019; Wiktorowicz et al., 2021).
4.3 Expectations for Antennae galaxy
The correlation between X-ray source positions and stellar clusters has been previously noted in starburst galaxies (e.g., Kaaret et al., 2004). Poutanen et al. (2013)’s study of the Antennae galaxies reveals a significant association between bright X-ray sources and stellar clusters, with most X-ray sources located outside clusters. This supports the idea that many ULXs are massive X-ray binaries ejected from their birth clusters, rather than IMBHs.
The displacements of ULXs from cluster centers, ranging up to $300$ parsecs, are statistically significant and likely due to ejection of massive binaries. The ejection mechanisms suggested by Poutanen et al. (2013), including supernova kicks and few-body encounters, are consistent with our results showing higher NS accretor fractions and smaller separations among escaped ULXs.
The distribution of ULXs likely evolves with stellar population age. While the Antennae study focus on young populations, older stellar systems are expected to decrease their X-ray binary populations (e.g., Fragos et al., 2008), highlighting the importance of considering population age in ULX distribution studies.
4.4 Future prospects
Future studies of ULXs in GCs should include more simulations to better estimate statistical effects, errors, and the significance of conclusions. A more consistent survey of the parameter space is needed, as well as the incorporation of evolutionary parameters (both stellar and binary evolution) beyond dynamical ones.
In the future work we plan to include the effect of beaming in ULXs (e.g., Lasota & King, 2023). Beaming increases the apparent luminosity of ULXs, but decreases the probability of observation due to possible misalignment (Wiktorowicz et al., 2019; Khan et al., 2022). Population synthesis simulations, Wiktorowicz et al. (e.g., 2019) indicate that the majority of NS ULXs are beamed. The beaming factor is dependent on the mass transfer rate, with higher mass transfer rates resulting in stronger beaming (King, 2009).
Additionally, the importance of wind Roche-lobe overflow (see e.g., Wiktorowicz et al., 2021; Zuo et al., 2021, and references therein) for NS ULXs with (super)giant donors should be explored. Wind Roche-lobe overflow can lead to stable mass transfer even with large mass ratios, and can boost mass transfer rates to reach ULX luminosity levels.
5 Conclusions
The apparent absence of ULXs in Milky Way GCs aligns more closely with our models of initially tidally filling clusters rather than non-tidally filling ones. This distinction is crucial: tidally filling clusters experience significant early mass loss through tidal stripping by their host galaxy’s gravitational field, leading to the ejection of potential ULX progenitors before they can evolve into ULXs. In contrast, non-tidally filling clusters maintain higher central densities for most of their life, allowing for enhanced ULX formation through dynamical interactions. The presence of IMBH ULXs exclusively in non-tidally filling simulations further emphasizes the importance of maintaining dense stellar environments for certain ULX formation channels.
Furthermore, our results suggest that field populations may be significantly ”polluted” by ULXs ejected from GCs (escapers), which could contribute to the relative underrepresentation of ULXs in GCs compared to field environments. Escapers can have properties very similar to the field populations or form distinct configurations.
Acknowledgements.
We are grateful to anonymous referee for their insightful comments which helped to improve the paper. GW, MG, AH, LH were supported by the Polish National Science Center (NCN) through the grant 2021/41/B/ST9/01191 AA acknowledges support for this paper from project No. 2021/43/P/ST9/03167 co-funded by the Polish National Science Center (NCN) and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339. For the purpose of Open Access, the authors have applied for a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
Data availability
The data underlying this article are available in Zenodo at https://doi.org/10.5281/zenodo.14953837. This includes simulation data on ULX systems, globular cluster properties, and initial binary conditions, along with corresponding header files containing column definitions.
Appendix A Additional tables
Table 7: Model parameters and initial properties
npop1-fb10-TF 1 10% $2.01$ $11.90$ $0.69$ $5.85$ $89.10$ $2.64$ $10,888.00$
npop1-fb10-nTF 1 10% $2.01$ $1.00$ $0.69$ $0.49$ $89.10$ $5.87$ $265.22$
npop1-fb95-TF 1 95% $1.96$ $11.93$ $0.72$ $5.81$ $93.93$ $2.61$ $6,386.70$
npop1-fb95-nTF 1 95% $1.96$ $1.00$ $0.72$ $0.49$ $93.93$ $5.84$ $155.06$
npop2-cpop05-fb95-nTF 2 0.05 95% $1.96$ $1.00$ $0.69$ $0.01$ $13.03$ $9.59$ $93.85$
npop2-cpop2-fb10-nTF 2 0.2 10% $2.01$ $1.00$ $0.66$ $0.04$ $12.84$ $7.77$ $174.58$
npop2-cpop2-fb95-nTF 2 0.2 95% $1.96$ $1.00$ $0.69$ $0.04$ $13.12$ $7.77$ $101.01$
131313 List of simulations with differentiating parameters and some initial properties (at $\mathrm{t}_{\mathrm{phys}}=0$). Label identifies particular simulation and will be used throughout the paper. TF/nTF in the label means initially tidally filling or non-tidally filling. NF, when present, means that new features were included in this model (see text for details); $\mathrm{N}_{\mathrm{pop}}$ number of stellar generations; $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$ ratio between the $\mathrm{R}_{\mathrm{h}}$ of the second and first stellar generation; $\mathrm{f}_{\mathrm{bin}}$ binary fraction; $\mathrm{R}_{\mathrm{gc}}$ galactocentric radius [$\mathrel{\mathrm{kpc}}$]; $\mathrm{R}_{\mathrm{h}}$ half-mass radius for the first population [$\mathrel{\mathrm{pc}}$]; $\mathrm{M}_{\mathrm{tot}}$ total stellar mass [$\mathrel{\mathrm{M}_{\odot}}$]; $\mathrm{R}_{\mathrm{c}}$ core radius [$\mathrel{\mathrm{pc}}$]; $\mathrm{M}_{\mathrm{c}}$ core mass [$\mathrel{\mathrm{M}_{\odot}}$]; $\rho_{\mathrm{c}}$ central density [$\mathrel{\mathrm{M}_{\odot}}/\mathrel{\mathrm{pc}}^{3}$]; $\mathrm{t}_{\mathrm{h,rel}}$ Spitzer half-mass relaxation time [$\mathrel{\mathrm{M}_{\odot}}$]. All simulations include the escapers information (see text). For all these simulations the counterpart field population (with dynamics turned off) was calculated (nodyn in the label; not shown in the table) with the same parameters.
@@ Appendix A Additional tables
141414 Fractional distribution of ULXs across different groups for various simulations. For group definitions see Table 2